OpenStudy (peterg2001):

help please will medal A 12-foot flagpole casts a nine-foot shadow. What is the measure from the top of the shadow to the top of the flag pole? 15 21 3 6 @SkyBudgie9303 @sky_guy02 @EclipsedStar

OpenStudy (anonymous):

i dunno. I'm SO sorry. im good at equations... not word problems..... but i'm logging off in 10 minutes so i wouldn't have been much help anyways... SORRY!

OpenStudy (anonymous):



OpenStudy (anonymous):

and the triangle equation...... a^2xB^2=c^2 right?

OpenStudy (anonymous):

plug the numbers in

OpenStudy (peterg2001):

answer? please?

OpenStudy (anonymous):

no answers tho..

OpenStudy (anonymous):

but i can give u the equation

OpenStudy (anonymous):

9^2+12^2=c^2

OpenStudy (anonymous):

81+144=c^2

OpenStudy (anonymous):

225=c^2

OpenStudy (anonymous):

and solve that